OpenStudy (anonymous):

On a certain map, a distance of 40 miles is represented by 1cm. How many miles are represented by 2.2cm on the map ?

OpenStudy (anonymous):

okay so we know that 2 cm is 80 miles, now we just need to find what the .2 on the map is

OpenStudy (anonymous):

Should I multiply ?

OpenStudy (anonymous):

okay, so if I recall, you should find 20% of 40 and add 80 to that, since the .2 in 2.2 is really 0.20 which is equal to 20%

OpenStudy (anonymous):

so add 80+8 I think

OpenStudy (anonymous):

Which is 160

OpenStudy (anonymous):

no

OpenStudy (anonymous):

80+8=88

OpenStudy (anonymous):

I mean 88

OpenStudy (anonymous):

right

OpenStudy (anonymous):

Sorry I'm not so focused

OpenStudy (anonymous):

I am not really sure if the .2 in 2.2 stands for a .02 or a .20

OpenStudy (anonymous):

but I think this is the answer: 88

OpenStudy (anonymous):

@OregonDuck

OpenStudy (mathstudent55):

Since a map scale is a ratio, map scale problems can usually be solved by a proportion.

OpenStudy (mathstudent55):

You are given a real distance and a map distance: Real distance Map distance 40 miles 1 cm This allows you to establish a ratio, which is the scale of the map: 40 miles : 1 cm

OpenStudy (mathstudent55):

Now you are given a map distance and an unknown real distance We can let x represent the unknown real distance. Real distance Map distance x 2.2 cm This can be written as a ratio: x : 2.2 cm

OpenStudy (mathstudent55):

The map scale ratio equals the ratio with the unknown real distance, so we write an equation showing the ratios are equal. An equation that is two equal ratios is called a proportion. Also, there are several ways of writing a ration. The ratio of x to y can be written in these ways: x to y x : y x/y Let's write an equation with our ratios, but at the same time let's change the ratios to the fraction form: \(\large \dfrac{40~miles}{1~cm} = \dfrac{x}{2.2~cm} \)

OpenStudy (anonymous):

Should I multiply 40 by 2.2 ?

OpenStudy (mathstudent55):

Now we need to solve the proportion for x. We cross multiply: \(1 \times x = 40 \times 2.2\) Now we simplify: \(x = 88\) Answer: the distance is 88 miles

OpenStudy (anonymous):

I understand it so much better now thank you so much I really appreciated that @mathstudent55 and @heretohelpalways :)

OpenStudy (anonymous):

I am glad to read/hear that. Happy Openstudying! :)

OpenStudy (mathstudent55):

You are very welcome. I'm glad you understand it better now.